OpenStudy (anonymous):

Given the function f(x) = 4(x + 1)2 − 3, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x−2) f(x) − 2 f(2x) 2•f(x)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

will medal need help !!!

OpenStudy (anonymous):

please help !!

OpenStudy (anonymous):

@jackunzel58

OpenStudy (anonymous):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

I'm not entirely sure. I am learning this right now so I'm not going to be much help :(

OpenStudy (anonymous):

do you know anyone that is good in math :(

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@ganeshie8

OpenStudy (haleyelizabeth2017):

@zpupster is good at math, I believe :)

OpenStudy (campbell_st):

so the function is \[f(x) =4(x + 1) ^2- 3\]

OpenStudy (anonymous):

yes @campbell_st

OpenStudy (anonymous):

@zpupster ??

OpenStudy (campbell_st):

ok... so do you know what the vertex is for the function...?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

ok... the vertex form is \[f(x) = a( x - h)^2 + k\] where (h, k) is the vertex. So comparing your function, what is the vertex of the parabola

OpenStudy (anonymous):

1 and -3

OpenStudy (campbell_st):

almost... (-1, -3) is the vertex f(x -2) means replace x with x - 2 so the function becomes \[f(x -2 = 4(x - 2 + 1)^2 - 3\] when you simplify this, does the vertex change... when you compare it to the original function..?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):